OpenStudy (anonymous):

trig integrals 1/sqrt (1-cos 2x)dx

OpenStudy (dumbcow):

1-cos(2x) = 1 - (cos^2 -sin^2) = (1-cos^2) +sin^2 = 2sin^2 \[\rightarrow \int\limits \frac{dx}{\sqrt{2} \sin x}\]

OpenStudy (anonymous):

and what would be the integral of 1/sqrt (sin 2x)